-4x+3y=6
y=4/3x-5
--- How many solutions does this system of linear equations have?

Answer by Alan3354(30993)

-4x+3y=6
y=4/3x-5 times 3 --> 3y = 4x - 15
--- How many solutions does this system of linear equations have?
----------
None